MILLIMETER AND SUBMILLIMETER WAVE SPECTRA OF THE HCOO^{13}CH_{3} ISOTOPOLOG OF METHYLFORMATE IN THE GROUND STATE AND IN THE FIRST EXCITED TORSIONAL STATE.

\underline{I. HAYKAL}, L. MARGULÈS, T. R. HUET, R. A. MOTIYENKO, Laboratoire PhLAM, UMR8523 CNRS-Université Lille 1, F-59655 Villeneuve d’Ascq Cedex, France; M. CARVAJAL, Unidad Asociada IEM-CSIC-Dpto. de Física Aplicada, Universidad de Huelva, Spain; I. KLEINER, LISA, CNRS et Université Paris Est et Paris Diderot, 61 av. Général de Gaulle, 94010, Créteil, France; J. C. GUILLEMIN, ENS de Chimie de Rennes -CNRS -35700 Rennes, France; B. TERCERO, J. CERNICHARO, Centro de Astrobiología (CSIC-INTA). Ctra de Ajalvir, Km 4, 28850 Torrejón de Ardoz, Madrid, Spain.

The detection of nineteen new rotational transitions of the parent molecule of methylformate (HCOOCH_{3}) in the second lowest excited torsional mode (ν_{t} = 2) was recently reported in Orion-KL^{a}, as well as the detection of eighty new lines corresponding to the two ^{18}O isotopologs of methylformate in their ground states^{b}. The laboratory work on HCOO^{13}CH_{3} was continued^{c}. A wide spectral range from 50 to 940 GHz was recorded in Lille with the submillimeter-wave spectrometer based on harmonic generation of a microwave synthesizer source, using a multiplication chain of solid state sources (50-100 and 150-940 GHz) and a backward wave oscillator (100-150 GHz), and coupled to a 2.2 m cell. The absolute accuracy of the line positions is better than 30 kHz up to 630 GHz and 50 kHz above. The two states (ν_{t} = 0 and 1) were fitted together using the RAM Hamiltonian of the BELGI program and a new set of 45 parameters was accurately determined. The fit contains 7050 lines corresponding to the ground state up to J = 78 and K_{a} = 34 and 1907 lines related to ν_{t} = 1 up to J = 59 and K_{a} = 24. The detection of new ν_{t} = 1 lines in Orion KL will be reported and discussed. This work is supported by the French Programme National de Physico-Chimie du Milieu Interstellaire (CNRS), by CNES, and by the Spanish Government through the grants FIS2011-28738-C02-02 and CONSOLIDER 2009-00038.

aS. Takano, Y. Sakai, S. Kakimoto, M. Sasaki, and K. Kobayashi PASJ. \underline{64}, 89, 2012.

bB. Tercero, et al. A& A. \underline{538}, A199, 2012.

cM. Carvajal, et al. A& A. \underline{500}, 1109, 2009.